Continue to assume that x = 5/4, but the market for long-distance telephone service remains perfectly competitive following the merger. For what values of s will the DOJ approve this merger?
1
Expert's answer
2020-06-04T10:22:22-0400

With the remaining stable indicators, we obtain the following expression for determining the quantitative characteristics of the quality of the services provided:


"Q=a+s-bp"


"s=Q-a+bp"


"s=x(K+\frac{1}{2}L)-a+bp"


"s=1.25(K+\frac{1}{2}L)-a+bp"